# Efficient Estimation of Unfactorizable Systematic Uncertainties

### v1

@@ 1 Introduction
In the following sections, we describe a strategy for characterizing the impact of systematic uncertainties without assuming factorization. Our strategy uses Gaussian processes (GPs) to regress observables of the detector response. Gaussian processes are flexible, nonparametric machine learning models that have been used in high-energy physics studies involving background selection and signal extraction [8, 9, 10]. While GPs have been less widely applied to the estimation of systematic uncertainties, we show that they have the potential to outperform more traditional approaches. In addition, we show that we can leverage derivative information when fitting the GPs by computing the gradient of the observables with respect to the nuisance parameters associated to the sources of uncertainty. We show that GPs that are fit with this additional derivative information can efficiently estimate the complex and unfactorizable effects of multiple sources of systematic uncertainty.
@@ 5.1 Gaussian Process Regression
Figure 2: Panel (a) shows the four training observations (red dots) used to fit the regular GP model. Also shown are the predicted mean $\mu_{\mathrm{test}}$ (blue line) and confidence band (light blue area), together with three functions sampled from the posterior (gray dotted lines). For reference, the true underlying function is depicted in pink. Similarly, panel (b) shows the predictions of the derivative GP model using the same four observations but augmented by their gradients (red tangent segments).
@@ 5.2 Bayesian Experimental Design
To evaluate the BED strategy, we perform 20 iterations. After each iteration, the predictive mean $(\mu_{\mathrm{test}})$ and the predictive covariance matrix $(\Sigma_{\mathrm{test}})$ of the models are updated. We evaluate the quality of the predictions by calculating the mean-squared-error (MSE), which provides a sense of accuracy. We also calculate the trace of the covariance matrix, $\mathrm{Tr}(\Sigma_{\mathrm{test}})$, which is proportional to the variance of the model, providing a sense of precision. Lower $\mathrm{Tr}(\Sigma_{\mathrm{test}})$ values suggest lower predictive uncertainty.
@@
Figure 3: First three BED iterations of the regular (a) derivative (b) GP models. The panels show the initial observations (red dots) and the observations selected by the BED strategy (orange dots). Also shown are the predictive mean $\mu_{\mathrm{test}}$ (blue line), the uncertainty band (light blue area), and three functions sampled from the posterior (gray dotted lines). For reference, the true underlying function is depicted in pink. Along the bottom of each panel, we show the utility function in green. The location of the next observation is indicated by the dotted vertical line, corresponding to the point that maximizes the utility.
@@ 6.2 Bayesian Experimental Design
Finally, we evaluate the use of the BED strategy by executing 44 iterations of the BED algorithm, leading to a total of 49 training samples, the same number used in the on-axis regression strategy. In particular, we examine the sensitivity of the BED strategy with respect to the number of grid points in the utility input $X_{u}$ and the regularization parameter $\gamma$. We consider the following scenarios: a $25\times 25$ grid with $\gamma^{2}=1/125$ and $\gamma^{2}=0$; a $10\times 10$ grid with $\gamma^{2}=1/50$ and $\gamma^{2}=0$; and a $5\times 5$ grid with $\gamma^{2}=1/10$ and $\gamma^{2}=0$. Two example grid configurations are shown in Fig. 9.
@@
7 High-Energy Physics Experiments: 4D MET Efficiency Estimation
@@ 8 Conclusions
Physics experiments, like those conducted at the Large Hadron Collider, play a crucial role in precision tests of the Standard Model, where experimental measures of particle properties are compared to theoretical predictions. A critical aspect of these tests is the accurate assessment of systematic uncertainties. Oftentimes, the significance of any observed discrepancy between measurement and theoretical prediction depends on correctly identifying all relevant sources of uncertainty and understanding their impact on the experimental response.
Ideally, after identifying the sources of uncertainty, one would conduct rigorous assessments to determine their effects on the experimental response. Such assessments would involve characterizing the space that the sources of uncertainty create. However, the complexity of this space scales exponentially with its dimensionality, making a thorough exploration and analysis a non-trivial task.
Several techniques have been employed to simplify the assessment of systematic uncertainties, which can be viewed as nuisance parameters. Most assume the factorization of their underlying correlations. These assumptions are often extended to treat the impact of the individual nuisance parameters on the experimental response as also factorizable variables. In this paper, we argue that such assumptions are not always reliable.
We propose a method for assessing the impact of systematic uncertainties in experiments without relying on the common assumption of factorization. Our approach uses Gaussian process regression to model observables of the experimental response as a function of the nuisance parameters. When analytical expressions for these observables are available, their gradients with respect to the nuisance parameters can be computed and incorporated into the regression. We show that GPs enhanced with derivative information can efficiently estimate the complex and unfactorizable effects of multiple sources of systematic uncertainty.
@@
K.C. is supported by NSF PHY-2323298. A.R. and D.W. are supported by the DOE Office of Science.

### v2

@@ 1 Introduction
In the following sections, we describe a strategy for characterizing the impact of systematic uncertainties without assuming factorization. Our strategy uses Gaussian processes (GPs) to regress observables of the detector response. Gaussian processes are flexible, nonparametric machine learning models that have been used in high-energy physics studies involving background selection and signal extraction [8, 9, 10]. While GPs have been less widely applied to the estimation of systematic uncertainties [11], we show that they have the potential to outperform more traditional approaches. In addition, we show that we can leverage derivative information when fitting the GPs by computing the gradient of the observables with respect to the nuisance parameters associated to the sources of uncertainty. We show that GPs that are fit with this additional derivative information can efficiently estimate the complex and unfactorizable effects of multiple sources of systematic uncertainty.
@@ 5.1 Gaussian Process Regression
Figure 2: Panel (a) shows the four training observations (red dots) used to fit the regular GP model. Also shown are the predicted mean $\mu_{\mathrm{test}}$ (blue line) and confidence band (light blue area), together with three functions sampled from the posterior (gray dotted lines). For reference, the true underlying function is depicted in magenta. Similarly, panel (b) shows the predictions of the derivative GP model using the same four observations but augmented by their gradients (red tangent segments).
@@ 5.2 Bayesian Experimental Design
We evaluate the BED strategy by performing 20 iterations of the algorithm. After each iteration, the predictive mean $(\mu_{\mathrm{test}})$ and the predictive covariance matrix $(\Sigma_{\mathrm{test}})$ of the models are updated. We evaluate the quality of the predictions by calculating the mean-squared-error (MSE), which provides a sense of accuracy. We also calculate the trace of the covariance matrix, $\mathrm{Tr}(\Sigma_{\mathrm{test}})$, which is proportional to the variance of the model, providing a sense of precision. Lower $\mathrm{Tr}(\Sigma_{\mathrm{test}})$ values suggest lower predictive uncertainty.
@@
Figure 3: First three BED iterations of the regular (a) derivative (b) GP models. The panels show the initial observations (red dots) and the observations selected by the BED strategy (orange dots). Also shown are the predictive mean $\mu_{\mathrm{test}}$ (blue line), the uncertainty band (light blue area), and three functions sampled from the posterior (gray dotted lines). For reference, the true underlying function is depicted in magenta. Along the bottom of each panel, we show the utility function in green. The location of the next observation is indicated by the dotted vertical line, corresponding to the point that maximizes the utility.
@@ 6.2 Bayesian Experimental Design
Finally, we evaluate the use of the BED strategy by executing 44 iterations of the BED algorithm, leading to a total of 49 training samples, the same number used in the on-axis regression strategy. In particular, we examine the sensitivity of the performance of the BED strategy with respect to the number of grid points in the utility input $X_{u}$ and the regularization parameter $\gamma$. We consider the following scenarios: a $25\times 25$ grid with $\gamma^{2}=1/125$ and $\gamma^{2}=0$; a $10\times 10$ grid with $\gamma^{2}=1/50$ and $\gamma^{2}=0$; and a $5\times 5$ grid with $\gamma^{2}=1/10$ and $\gamma^{2}=0$. Two example grid configurations are shown in Fig. 9.
@@
7 High-Energy Physics Experiments: 4D Efficiency Estimation
@@ 8 Conclusions
Physics experiments, like those conducted at the Large Hadron Collider, play a crucial role in precision tests of the Standard Model, where experimental measures of particle properties are compared to theoretical predictions. A critical aspect of these tests is the accurate assessment of systematic uncertainties.
Ideally, after identifying the sources of uncertainty that affect an analysis, one would thoroughly and quantitatively map out the effects of those uncertainties on the experimental response. It is typically straightforward to associate nuisance parameters to the sources of uncertainty; however, estimating the response as a function of the nuisance parameters is challenging – it corresponds to estimating a function of several variables and is subject to the curse of dimensionality.
Several techniques have been introduced to estimate the effect of multiple sources of systematic uncertainties. Most assume the individual sources of uncertainty are independent, though it is fairly straightforward to incorporate correlations. More critically, most approaches also assume that the effect of those uncertainties factorizes, which is much less motivated. In this paper, we provide examples where these assumptions fail and propose a method for estimating the joint response to multiple sources of systematic uncertainties without assuming they factorize. Our approach uses Gaussian process regression to model observables of the experimental response as a function of the nuisance parameters. When (approximate) gradients of the response with respect to the nuisance parameters can be computed, that informaiton can also be incorporated into the regression. We show that GPs enhanced with derivative information can efficiently estimate the complex and unfactorizable effects of multiple sources of systematic uncertainty.
@@
K.C. was supported by NSF PHY-2323298. A.R. and D.W. are supported by the DOE Office of Science.